### v4

# Neural Dynamic Data Valuation: A Stochastic Optimal Control Approach

Abstract
Data valuation has become a cornerstone of the modern data economy, where datasets function as tradable intellectual assets that drive decision-making, model training, and market transactions. Despite substantial progress, existing valuation methods remain limited by high computational cost, weak fairness guarantees, and poor interpretability, which hinder their deployment in large-scale, high-stakes applications. This paper introduces Neural Dynamic Data Valuation (NDDV), a new framework that formulates data valuation as a stochastic optimal control problem to capture the dynamic evolution of data utility over time. Unlike static combinatorial approaches, NDDV models data interactions through continuous trajectories that reflect both individual and collective learning dynamics. The framework integrates a meta-learned re-weighting mechanism based on mean-field interactions to ensure fair value assignment across heterogeneous samples and employs the interpretable Kolmogorov–Arnold Networks (KANs) with Matérn kernels to reveal how each data point contributes to model outcomes. Extensive experiments on six real-world datasets demonstrate that NDDV achieves up to 58× computational speed-up, improves fairness metrics by a large margin, and consistently yields higher F1-scores for corrupted-data detection and utility estimation compared with state-of-the-art methods. By unifying efficiency, fairness, and interpretability under a single stochastic-control formulation, NDDV offers a scalable and transparent paradigm for trustworthy data valuation in practical machine learning systems.
Index Terms:
Data economy, Data valuation, Marginal contribution, Stochastic optimal control
1 Introduction
Data has become a central asset in the modern economy, functioning as intellectual property that drives value creation in data marketplaces [1, 2, 3]. The intrinsic worth of data facilitates sharing, exchange, and reuse across organizations, businesses, and research communities. This value arises from multiple intertwined factors such as data quality, market demand, computational sophistication of downstream tasks, and data integrity, all of which determine the reliability and utility of data across domains.
In the current data ecosystems, the capacity to measure and assign equitable value to data is essential for maintaining trust, transparency, and sustainability in data exchange. Accurate valuation enables fair compensation for data providers and informed acquisition decisions for consumers, allowing data resources to be allocated efficiently. Moreover, in machine learning and artificial intelligence, the performance of predictive models depends strongly on the contribution of individual data points to overall model utility. Quantifying these contributions has therefore become a foundational mechanism for model accountability, data quality assurance, and the efficient operation of data marketplaces. As data volume and diversity continue to grow rapidly, developing systematic, interpretable, and efficient data valuation mechanisms has become indispensable for the sustainable development of trustworthy data-driven systems.
The development of an effective data valuation framework encounters several intertwined technical difficulties that arise from the intrinsic characteristics of large-scale learning systems. The first difficulty lies in computational scalability. Estimating the individual contribution of each data point requires measuring subtle variations in model utility under high-dimensional and dynamically changing training processes. As data volume grows, even modest estimation errors can propagate rapidly, making valuation unstable or infeasible without a computationally efficient mechanism. A second difficulty concerns fairness in contribution assessment. Data points differ widely in quality, representativeness, and influence on learning outcomes. Accurately reflecting these heterogeneous effects while preventing bias amplification demands a formulation that adapts to the evolving collective behaviour of data rather than applying uniform weighting. The third difficulty involves interpretability of valuation dynamics. Because data value evolves throughout model training, a transparent framework must reveal how and why these values change across layers and epochs. Providing such temporal traceability is essential for trustworthy decision support and reliable auditing of data assets. Overcoming these difficulties requires a unified theoretical model capable of handling dynamic interactions, ensuring equitable valuation, and explaining the evolution of data influence in an efficient and interpretable manner.
Earlier research on data valuation has evolved from model-specific influence estimation to cooperative game formulations and recent efficiency-driven approximations. The influence-based approach [4] estimates each sample’s importance through sensitivity analysis of model parameters but struggles to capture complex dependencies among data points. The game-theoretic framework [5, 6, 7, 8, 9] assigns value through average marginal contributions, offering strong fairness principles but requiring exponentially many evaluations. To mitigate this burden, efficient approximation techniques have been introduced, such as neighbor-based models [10, 11, 12] and ensemble reweighting strategies [13, 14], yet these often compromise fairness or interpretability. Further, explainability-oriented studies [15] enhance transparency in local model reasoning but do not extend to global data valuation. Despite these advances, existing methods remain limited by three unresolved issues: excessive computational cost that restricts scalability, biased or inconsistent value allocation across heterogeneous data, and lack of transparency regarding how value evolves during learning. These unresolved problems motivate a unified framework that models data value dynamically, enabling efficient, fair, and interpretable estimation at scale [1, 2, 3].
To overcome these limitations, we propose Neural Dynamic Data Valuation (NDDV), which redefines data valuation as a stochastic optimal control problem. This novel perspective models data value as a time-evolving trajectory rather than a static quantity, allowing the framework to capture both individual and collective learning dynamics. The first module formulates the valuation process through forward–backward stochastic differential equations, providing a unified trajectory-based learning mechanism that drastically improves efficiency. The second module introduces a fairness-aware re-weighting mechanism grounded in mean-field interactions, which adaptively balances individual and group contributions to mitigate bias. The third module enhances interpretability by adopting Kolmogorov–Arnold Networks (KANs) with Matérn kernels, which yield transparent mappings between data points and their utility trajectories. Together, these components form a coherent system that achieves scalability, fairness, and transparency simultaneously—three goals that prior methods could not integrate effectively.
The key contributions of this paper are summarized as follows:
  • Unified Stochastic Control Formulation. We reformulate data valuation as a stochastic optimal control problem, allowing the dynamic modeling of data-value trajectories through a single training process, which achieves substantial efficiency improvement over traditional retraining-based methods.
  • Fairness-Aware Mean-Field Reweighting. We propose a meta-learned re-weighting strategy that integrates individual heterogeneity and collective-state interactions, ensuring equitable valuation and reducing bias amplification across data subgroups.
  • Interpretable Utility Representation. We design a KANs architecture with Matérn kernels that reveals how data contributions evolve over layers and epochs, thereby improving interpretability and auditability.
  • Comprehensive Empirical Validation. Extensive experiments on six benchmark datasets demonstrate that NDDV achieves up to 58× speed-up, higher fairness scores, and superior corrupted-data detection performance compared with state-of-the-art methods.
The remainder of this paper is structured as follows. Section 2 reviews existing research on data valuation and optimal control theory. Section 3 introduces the preliminaries and presents the mathematical formulation of the problem. Section 4 details the NDDV framework, including the stochastic control formulation, fairness-aware reweighting, and interpretable architecture. Section 5 provides theoretical and complexity analyses. Section 6 presents experimental settings, results, and discussions. Section 7 concludes the paper and outlines future research directions.
Figure 1: Neural dynamic data valuation schematic and results. The panel compares NDDV and existing methods. It is evident that NDDV transforms the static combined evaluation method of existing data valuation into a dynamic optimization process, defining a new utility function and dynamic marginal contribution. Compared to existing methods, NDDV requires only one training session to determine the value of all data points, significantly enhancing computational efficiency. Taking the half-moon dataset as an example, we demonstrate some results of NDDV to indicate its effectiveness.
2 Related works
2.1 Dynamics and Optimal Control Theory
The dynamical interpretation of deep neural networks (DNNs) as continuous-time systems has introduced new theoretical perspectives on learning and optimization. Foundational studies view deep residual architectures as discrete approximations of ordinary differential equations (ODEs), bridging neural network propagation and numerical integration theory [16, 17, 18]. This connection enables analysis in the continuum limit through Wasserstein geometry, which links gradient flow and optimal transport [19]. The resulting framework naturally extends to mean-field control formulations for learning dynamics [20, 21].
Building on this perspective, several optimization algorithms have been reformulated as control problems [22, 23]. Stochastic gradient descent (SGD) has been analyzed through the lens of stochastic dynamical systems, where Langevin and Lévy dynamics describe implicit regularization effects [24, 25]. Further, studies of Gram matrix evolution have demonstrated convergence and global optimality properties [26, 27], providing guidance for adaptive tuning of hyperparameters such as learning rate and batch size [28, 29]. Collectively, these results suggest that the theory of dynamic systems and optimal control offers a principled foundation for reinterpreting neural learning and motivates the control-based formulation adopted in this work.
2.2 Data Valuation
Data valuation has emerged as a key research topic to quantify each sample’s contribution to model performance. Early approaches such as Leave-One-Out (LOO) [4] and Data Shapley [10, 5] estimate marginal contributions by iteratively removing data points and observing model performance changes. Although theoretically grounded, these methods suffer from exponential computational cost. Extensions like Beta Shapley [9] and Data Banzhaf [30] improve fairness modeling but remain computationally demanding. Influence-based estimators [31] leverage gradient approximations to reduce complexity, yet often lose robustness in nonconvex learning settings.
To enhance scalability, proxy and ensemble-based techniques were introduced. KNNShapley [11] employs a $k$-Nearest Neighbor classifier as a learning-agnostic surrogate, while AME [13] and Data-OOB [14] adopt ensemble reweighting and out-of-bag estimators for efficiency. These methods, however, still struggle to maintain fairness and transparency. Algorithm-agnostic formulations, such as volume-based valuation [32] and Wasserstein-distance metrics [33], bypass model retraining but lack the ability to detect label errors or explain valuation outcomes.
Beyond standard utility estimation, marginal contribution analysis has been extended to feature attribution [34, 35], interpretability [36, 37], and collaborative learning [38, 39]. Further refinements [40, 41] relax classical Shapley axioms to improve efficiency, while reinforcement learning-based estimators [42] model data usage likelihood during training. Despite these developments, current frameworks still face three inherent constraints: high computational cost, unfair valuation under data heterogeneity, and limited interpretability.
To our knowledge, optimal control has not yet been exploited for data valuation problems. Compared with the existing work, the proposed Neural Dynamic Data Valuation (NDDV) introduces a dynamic optimal control formulation that transforms valuation from static combinatorial computation to continuous-time learning dynamics. By coupling data points with a mean-field state, NDDV achieves efficient one-pass value estimation, fairness-aware reweighting, and interpretable trajectory analysis, which effectively addresses the key deficiencies observed in existing approaches.
3 Preliminaries and Problem formulation
Definition 1 (Leave-One-Out (LOO) Metric).
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ denote a training dataset, and let $U:2^{[N]}\rightarrow\mathbb{R}$ be a utility function that maps any subset of indices $S\subseteq[N]$ to a real-valued performance score. The LOO value of a data point $(x_{i},y_{i})\in\mathcal{D}$ is defined as:
$\phi_{\mathrm{loo}}(x_{i},y_{i};U)\triangleq U([N])-U([N]\setminus\{i\}).$ (1)
which quantifies the change in utility when the $i$-th point is excluded from the full dataset. This metric approximates the influence of a data point but can yield inaccurate values close to zero in practice [43].
Definition 2 (Static Marginal Contribution).
Let $(x_{i},y_{i})\in\mathcal{D}$ and $j\in[N]$. The marginal contribution of point $(x_{i},y_{i})$ with respect to subsets of size j1j-1 is defined as:
$\Delta_{j}(x_{i},y_{i};U)\triangleq\frac{1}{\binom{N-1}{j-1}}\sum_{S\in\mathcal{D}^{\setminus(x_{i},y_{i})}_{j}}\left[U(S\cup\{i\})-U(S)\right].$ (2)
where $\mathcal{D}^{\setminus(x_{i},y_{i})}_{j}=\left\{S\subseteq[N]\setminus\{i\}:|S|=j-1\right\}$ is the set of all subsets of $\mathcal{D}\setminus\{(x_{i},y_{i})\}$ of size j1j-1. This static measure estimates the utility gained by adding $(x_{i},y_{i})$ into subsets of a fixed size.
Definition 3 (Shapley Value).
The Shapley value of a data point $(x_{i},y_{i})\in\mathcal{D}$ is defined as the average of its marginal contributions across all subset sizes:
$\phi_{\mathrm{Shap}}(x_{i},y_{i};U)\triangleq\frac{1}{N}\sum_{j=1}^{N}\Delta_{j}(x_{i},y_{i};U).$ (3)
This metric assigns values based on the average utility improvement when the data point is added to subsets of varying sizes.
3.1 Problem Formulation
Data valuation aims to quantify the importance of each training instance in determining the performance of a machine learning model. Consider a supervised learning setting with a training dataset $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$, where each $x_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$ denotes a data instance and $y_{i}\in\mathcal{Y}\subset\mathbb{R}$ is the corresponding label. Let $N]:={1,2,\dots,N}$ index the data points in $\mathcal{D}$.
We define a utility function $U:2^{[N]}\rightarrow\mathbb{R}$, which maps any subset $S\subseteq[N]$ to a scalar performance score obtained by training a learning algorithm on the subset $\mathcal{D}S={(x_{i},y_{i})}{i\in S}$ and evaluating the resulting model on a fixed validation set. The function $U$ serves as a surrogate for model quality induced by data subset $S$, and $U(\emptyset)$ is defined as the performance of a reference model trained without any data, typically a constant predictor.
The objective of data valuation is to assign a value $\phi_{i}\in\mathbb{R}$ to each data point $(x_{i},y_{i})$ such that the collection ${\phi_{i}}_{i=1}^{N}$ reflects the individual contribution of each point to the overall utility of the dataset. That is, the mapping
$\phi:\mathcal{D}\rightarrow\mathbb{R},\quad(x_{i},y_{i})\mapsto\phi_{i}=\phi(x_{i},y_{i};U)$
quantifies how essential or beneficial each training point is to model performance.
4 Method
The proposed Neural Dynamic Data Valuation (NDDV) framework reformulates the data valuation problem as a stochastic optimal control process, where data contribution evolves continuously within the learning dynamics of a neural network. NDDV comprises three tightly integrated modules that operate coherently to achieve these objectives. The first module, the Dynamic State Encoder, transforms data samples into a latent representation governed by stochastic differential equations, which model their dynamic interactions with the evolving learning system [24, 23]. The second module, the Fairness-Aware Mean-Field Controller, defines a mean-field potential that regulates global fairness across samples by adjusting their control weights according to their statistical influence, ensuring equitable value allocation even under data heterogeneity [44, 45]. The third module, the Interpretable Neural Valuator, leverages KANs with Matérn kernels [46, 47] to map the evolving latent states into valuation scores, preserving the underlying dynamics while offering transparency in the contribution estimation. Collectively, these modules constitute a unified control-driven valuation framework that jointly ensures efficiency, fairness and interpretability. This design provides a principled and scalable solution to overcome the core deficiencies of prior valuation paradigms and enables reliable deployment of data valuation in large-scale machine learning systems.
We provide a pseudo algorithm in Alg.1 to illustrate the process of NDDV. It is evident that the implementation of NDDV is straightforward and easy to implement.
Algorithm 1 Pseudo-code of NDDV training
1:Training data $\mathcal{D}$, meta-data set $\mathcal{D}^{{}^{\prime}}$, batch size $n,m$, max iterations $K$.
2:The value of data points: $\phi(x_{i},y_{i};U(S))$.
3:Initialize The base optimization parameter $\psi^{0}$ and the meta optimization parameter $\theta^{0}$.
4:for $k=0$ to K1K-1 do
5:  $\{x,y\}\leftarrow$ SampleMiniBatch($\mathcal{D},n$).
6:  $\{x^{{}^{\prime}},y^{{}^{\prime}}\}\leftarrow$ SampleMiniBatch($\mathcal{D}^{{}^{\prime}},m$).
7:  Formulate the base training function $\hat{\psi}^{k}(\theta)$ by
$\hat{\psi}_{t}^{k}=\psi_{t}^{k}+\frac{\alpha}{N}\sum\limits_{i=1}^{N}\sum\limits_{t=0}^{T-1}\nabla_{\psi}\mathcal{H}_{i}(\cdot,\psi_{t}^{k},\mathcal{V}({\Phi_{i}(\cdot,\psi^{k}_{T});\theta})){\big|_{\psi^{k}}},$
8:  Update the base optimization parameters $\theta^{k+1}$ by
$\theta^{k+1}=\theta^{k}-\frac{\beta}{M}\sum\limits_{i=1}^{M}\nabla_{\theta}\ell_{i}(\hat{\psi}^{k}(\theta)){\big|_{\theta}^{k}},$
9:  Update the meta optimization parameters $\psi^{k+1}$ by
$\psi_{t}^{k+1}=\psi_{t}^{k}+\frac{\alpha}{N}\sum\limits_{i=1}^{N}\sum\limits_{t=0}^{T-1}\nabla_{\psi}\mathcal{H}_{i}(\vbox{\hbox{.}},\psi_{t}^{k},\mathcal{V}({\Phi_{i}(\cdot,\psi^{k}_{T});\theta^{k+1}})){\big|_{\psi^{k}}},$
10:  Update the weighted mean-field state $\mu^{k+1}_{t}$ by
$\mu^{k+1}_{t}=\frac{1}{N}\sum_{i=1}^{N}\sum\limits_{t=0}^{T-1}\mathcal{V}({\cdot,\mu^{k+1}_{i,T});\theta^{k+1}})X^{k+1}_{i,t}.$
11:end for
12:Compute the data state utility function $U_{i}(S)$ by
$U_{i}(S)=-X_{i,T}\cdot Y_{i,T},$
13:Compute the dynamic marginal contribution $\Delta(x_{i},y_{i};U_{i})$ by
$\Delta(x_{i},y_{i};U_{i}(S))=U_{i}(S)-\sum_{j\in\{1,\ldots,N\}\setminus i}\frac{U_{j}(S)}{N-1},$
14:Compute the value of data points $\phi(x_{i},y_{i};U(S))$ by
$\phi(x_{i},y_{i};U_{i}(S))=\Delta(x_{i},y_{i};U_{i}(S)).$
4.1 Learning Dynamic Valuation Trajectories via Stochastic Control
To address the inefficiency and static limitations of conventional data valuation, we formulate valuation as a stochastic optimal control problem [48, 49]. This dynamic view models how data points jointly influence learning over time, where each point follows a stochastic trajectory capturing its evolving contribution (see Fig. 2).
Let $\psi:[0,T]\rightarrow\Psi\subset\mathbb{R}^{p}$ denote the control parameters and $X_{t}=(X_{1,t},\dots,X_{N,t})\in\mathbb{R}^{d\times N}$ the data-state vector. The expected cost functional is defined as:
$\mathcal{L}(\psi)=\mathbb{E}\!\left[\int_{0}^{T}R(X_{t},\psi_{t})\,\mathrm{d}t+\Phi(X_{T},\psi_{T})\right],$ (4)
where $R$ and $\Phi$ are the running and terminal costs, respectively.
The state evolves according to a forward stochastic differential equation (FSDE):
$\mathrm{d}X_{t}=b(X_{t},\psi_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t},\qquad X_{0}=x,$ (5)
where $b$ is the drift function, $\sigma$ is the diffusion coefficient, and $W_{t}$ is a standard Wiener process.
Applying the Stochastic Maximum Principle (SMP) [50], we introduce the Hamiltonian:
$\mathcal{H}(X_{t},Y_{t},Z_{t},\psi_{t})=b(X_{t},\psi_{t})\!\cdot\!Y_{t}+\mathrm{tr}(\sigma^{\top}Z_{t})-R(X_{t},\psi_{t}),$ (6)
and the backward stochastic differential equation (BSDE) for the co-state $Y_{t}$:
$\mathrm{d}Y_{t}=-\nabla_{x}\mathcal{H}(X_{t},Y_{t},\psi_{t})\,\mathrm{d}t+Z_{t}\,\mathrm{d}W_{t},\qquad Y_{T}=-\nabla_{x}\Phi(X_{T}).$ (7)
The coupled FSDE–BSDE system forms a McKean–Vlasov FBSDE [51, 44, 52], and the optimal control $\psi_{t}^{*}$ satisfies:
$\mathcal{H}(X_{t}^{*},Y_{t}^{*},Z_{t}^{*},\psi_{t}^{*})\geq\mathcal{H}(X_{t}^{*},Y_{t}^{*},Z_{t}^{*},\psi),\quad\forall\,\psi.$ (8)
Defining Data Utility. For each data point $(x_{i},y_{i})$, its terminal utility is
$U_{i}(S)$ $=X_{i,T}\!\cdot\!\frac{\partial\mathcal{L}(\psi)}{\partial X_{i,T}}$
$=X_{i,T}\!\cdot\!\nabla_{x}\Phi_{i}(X_{i,T},\mu_{T},\psi_{i,T})$
$=-X_{i,T}\!\cdot\!Y_{i,T},$ (9)
linking valuation to the gradient of the cost function [53, 54].
The marginal and dynamic valuation score is defined as:
$\phi(x_{i},y_{i};U_{i}(S))=U_{i}(S)-\frac{1}{N-1}\sum_{j\neq i}U_{j}(S),$ (10)
indicating whether a sample benefits ($\phi>0$) or harms ($\phi<0$) learning.
Temporal Valuation Dynamics. The joint evolution of $X_{t}$ and $Y_{t}$ reveals how each sample’s influence changes across layers and epochs. The layer-epoch valuation is given by:
$\phi^{k}_{i,t}=-X^{k}_{i,t}\!\cdot\!Y^{k}_{i,t}+\sum_{j\neq i}\frac{X^{k}_{j,t}\!\cdot\!Y^{k}_{j,t}}{N-1},$ (11)
allowing continuous tracking of sample importance during training.
This stochastic dynamic formulation eliminates combinatorial retraining, models temporal evolution of data influence, and offers interpretability across layers and epochs, the capabilities absent in static valuation heuristics [55, 56].
Figure 2: Learning data stochastic dynamic schematic. a. In stochastic optimal control, data points get their optimal state trajectories via dynamic interactions with the mean-field state. b. Within the data re-weighting strategy, data points are characterized by heterogeneity. In this scenario, data points dynamically interact with the weighted mean-field state, thereby determining their optimal state trajectories.
4.2 Fairness-Aware Valuation via Reweighting and Mean-Field Dynamics
To ensure fairness in data valuation, NDDV integrates adaptive reweighting and mean-field control with formal fairness guarantees. This mechanism adjusts both individual and group-level influences, addressing biases that arise under static uniform treatment.
Adaptive Reweighting via Meta-Learned Weights. Each data point is dynamically reweighted by a meta-learned function $V(\Phi_{i}(X_{T});\theta)$ inspired by Stackelberg game theory [45]. The control objective becomes:
$L(\psi;\theta)=\frac{1}{N}\sum_{i=1}^{N}\!\left[\sum_{t=0}^{T-1}R_{i}(\cdot,\psi_{t})+V(\Phi_{i}(\cdot,\psi_{T});\theta)\Phi_{i}(\cdot,\psi_{T})\right],$ (12)
where $\psi$ is the control path and $\Phi_{i}$ denotes terminal loss. The meta-network uses ReLU–Sigmoid activations to constrain $V\!\in[0,1]$.
The reweighting parameters $\theta$ are optimized via a bi-level problem:
$\begin{cases}\psi^{*}(\theta)=\arg\min_{\psi}L(\psi;\theta),\\ \theta^{*}=\arg\min_{\theta}\frac{1}{M}\sum_{i=1}^{M}\ell_{i}(\psi^{*}(\theta)),\end{cases}$ (13)
solved efficiently through a single-loop online update [57, 58]. This procedure amplifies representative and fair samples while reducing bias from noisy or minority points.
Mean-Field Interaction and Weighted Aggregation. Group fairness is enforced by embedding reweighted data into a mean-field dynamic system [44]:
$\mathrm{d}X_{i,t}=b(X_{i,t},\mu_{t},\psi_{i,t})\,\mathrm{d}t+\sigma\,\mathrm{d}W_{i,t},$ (14)
where the weighted population state is
$\mu_{t}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{V}(\Phi_{i}(\cdot,\psi_{i,T});\theta)X_{i,t}.$ (15)
With a linear–quadratic drift [59, 45, 60],
$\mathrm{d}X_{t}=[a(\mu_{t}-X_{t})+\psi_{t}]\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t},$ (16)
the discrete optimization becomes
$\min_{\psi,\theta}\frac{1}{N}\!\sum_{i=1}^{N}\!\!\Big[\!\sum_{t=0}^{T-1}\!R_{i}(\cdot,\psi_{i,t})+\mathcal{V}(\Phi_{i}(\cdot,\psi_{i,T});\theta)\Phi_{i}(\cdot,\psi_{i,T})\!\Big],$
$\text{s.t.}\;X_{i,t+1}=X_{i,t}+[a(\mu_{t}-X_{i,t})+\psi_{i,t}]\Delta t+\sigma\Delta W.$ (17)
The associated fairness-aware Hamiltonian is
$\mathcal{H}_{i}(\cdot,\psi_{t})=[a(\mu_{t}-X_{i,t})+\psi_{t}]\!\cdot\!Y_{i,t}+\sigma^{\top}Z_{i,t}-R_{i}(\cdot,\psi_{t}).$ (18)
Formal Fairness Guarantee. Following [5, 61], NDDV satisfies symmetry, monotonicity, and null-player properties. Fairness is quantified by the Equal Opportunity (EOp) and Equalized Odds (EOdds) scores [62]:
$\begin{cases}\text{EOp}_{i}=|\text{TPR}-\text{TPR}_{i}|,\\ \text{EOdds}_{i}=|\text{TPR}-\text{TPR}_{i}|+|\text{FPR}-\text{FPR}_{i}|.\end{cases}$ (19)
Lemma 1 (Bounded Fairness Violation under Meta-Reweighting).
Let $\mathcal{V}(\cdot;\theta)$ be Lipschitz-continuous with range $[0,1]$, and let dynamics follow Eq. (16). Then, for any subgroup $G_{i}$,
$\big|\mathbb{E}_{G_{i}}[\Phi]-\mathbb{E}_{\mathcal{D}}[\Phi]\big|\leq C\,\Delta_{\mathcal{V}},$ (20)
where $\Delta_{\mathcal{V}}=\max_{(x,y)}\mathcal{V}(\Phi(x,y);\theta)-\min_{(x,y)}\mathcal{V}(\Phi(x,y);\theta)$ and $C$ depends on the Lipschitz constant of $\Phi$.
Proof sketch.
By decomposing group-wise loss differences and bounding with the range of $\mathcal{V}$, the disparity scales with $\Delta_{\mathcal{V}}$. As $\mathcal{V}$ is smooth and bounded, fairness remains controlled across groups. ∎
This bound ensures that valuation disparities are limited when meta-weights remain stable. NDDV thus provides a principled, efficient, and fairness-aware approach to data valuation.
Figure 3: Data value trajectories at the layer-wise level.
Figure 4: Data value trajectories at the epoch-wise level.
Figure 5: Revealing the data valuation process. Using the 2dplanes dataset as an example, we assessed the impact on test accuracy via the removal and addition of high/low-value trajectories.
4.3 Learning Interpretable Utility Function
To enhance transparency in data valuation, NDDV learns an interpretable utility function by explicitly modeling the factors that shape data dynamics. The drift term $b(X_{t},\mu_{t},\psi_{t})$, which governs data evolution, depends on both the control policy $\psi$ and the reweighting parameters $\mathcal{V}$.
Inspired by KANs [46], we replace dense weight matrices with univariate basis functions, expressing $\psi$ and $\mathcal{V}$ as compositions of nonlinear mappings for improved interpretability and analytical control. The control strategy $\psi$ is represented as:
$\psi(X_{t})$ $=\sum_{q=1}^{2d+1}H_{t,q}\!\left(\sum_{p=1}^{d}h_{q,p}(X_{t})\right)$
$=\left(H_{T-1}\!\circ\!H_{T-2}\!\circ\!\cdots\!\circ\!H_{1}\!\circ\!H_{0}\right)\!X_{t},$ (21)
where $h_{q,p}$ are learnable univariate activations, and $H_{t}$ denotes transformation matrices derived from their superpositions.
Each activation is defined through a residual combination:
$h_{q,p}(X_{t})=\alpha_{b}h_{b}(X_{t})+\alpha_{k}h_{k}(X_{t},c),$ (22)
where $h_{b}$ uses the SiLU function, and $h_{k}$ is a radial basis function (RBF) centered at $c$. Coefficients $\alpha_{b}$ and $\alpha_{k}$ are initialized via Xavier and optimized during training.
Because Gaussian RBFs can over-smooth noisy distributions, we adopt a **Matérn kernel** [47] to improve local adaptivity:
$h_{k}(X_{t},c)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\,|X_{t}-c|}{\ell}\right)^{\!\nu}K_{\nu}\!\left(\frac{\sqrt{2\nu}\,|X_{t}-c|}{\ell}\right),$
where $\ell$ is the length scale, $\nu>0$ controls smoothness, $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ the modified Bessel function. As $\nu\!\to\!\infty$, this kernel approaches a Gaussian. For $\nu\!=\!3/2$, we obtain:
$h_{k}^{3/2}(X_{t},c)=\!\left(1+\frac{\sqrt{3}|X_{t}-c|}{\ell}\right)\exp\!\left(-\frac{\sqrt{3}|X_{t}-c|}{\ell}\right).$
Conventional neural valuation networks obscure how data points influence control dynamics. KANs with Matérn kernels decompose transformations into interpretable univariate components, enabling inspection of learned effects over time. This structure reveals how each data point’s contribution evolves, supporting transparent auditing and debugging of valuation behavior.
5 Computational Complexity Analysis
We analyze the computational complexity of the NDDV framework by decomposing it into its main components. Let $n$ denote the number of training data points, $b$ the batch size used in model training, and $k$ the number of meta-training iterations. Assume that forward and backward passes through the model take constant time per batch.
  1. 1.
    Meta-Gradient Training: NDDV leverages meta-gradient optimization for learning utility functions. For each iteration, a batch of size $b$ is used. Across $k$ iterations, the total cost is: $\mathcal{O}\left(\frac{kn}{b}\right)$, since all $n$ points are processed multiple times in mini-batches.
  2. 2.
    Forward-Backward Dynamics Learning: The stochastic control-based learner uses a forward-backward process across $T$ time layers, but each layer only propagates through a subset of data. Since $T$ is constant and independent of $n$, the total cost is dominated by the meta-training step and remains: $\mathcal{O}\left(\frac{kn}{b}\right)$.
  3. 3.
    Utility Evaluation: Once trained, the model produces utility scores directly without requiring individual retraining or retracing for each data point, as in leave-one-out methods. Thus, utility assignment incurs negligible additional cost.
Overall Complexity. Combining the steps above, the overall computational complexity of NDDV is: $\mathcal{O}\left(\frac{kn}{b}\right)$.
Comparative Efficiency. For reference, the complexities of competing methods are:
  • KNNShapley: $\mathcal{O}(n^{2}\log n)$, due to pairwise distance computations and sorting.
  • AME: $\mathcal{O}(kn/b)$, similar to NDDV but without forward-backward modeling.
  • Data-OOB: $\mathcal{O}(Bdn\log n)$, where $B$ is the number of trees and $d$ is the number of features.
Implications. Unlike methods that scale quadratically or depend on ensemble size and feature dimension, NDDV achieves linear complexity in $n$. This makes it highly scalable for large datasets while preserving competitive performance. The batch-wise meta-learning structure further reduces memory footprint and supports parallelization.
6 Experiments
6.1 Experimental Setup
Datasets. We evaluate our approach using six publicly available datasets provided in OpenDataVal [63], several of which have been widely adopted in prior work [5, 9]. These datasets span diverse domains—including image, text, and tabular data with varying degrees of class imbalance, feature sparsity, and task complexity, as detailed in Table I. This diversity ensures a comprehensive assessment of both the scalability and generalizability of our method. Moreover, the inclusion of datasets previously used in influential studies facilitates direct comparisons and fair benchmarking.
Baselines. We compare NDDV against eight representative data valuation methods: LOO [4], DataShapley [5], BetaShapley [9], DataBanzhaf [30], InfluenceFunction [31], KNNShapley [11], AME [13], and Data-OOB [14]. These methods cover a wide spectrum of valuation paradigms, including exact and approximate game-theoretic methods, influence-based techniques, and ensemble-based approaches. This selection ensures a balanced evaluation across accuracy, efficiency, and robustness. For fairness, all baselines are run with the same or greater number of utility evaluations than NDDV.
Experimental Setup. All experiments are conducted using a Python-based implementation on a machine equipped with an Intel Xeon Gold 6126 CPU @ 2.60GHz, 256GB RAM, and an NVIDIA Tesla V100 GPU with 32GB memory. The software environment includes Python 3.10, PyTorch 2.0, and CUDA 11.7. Our implementation is publicly available at https://github.com/liangzhangyong/NDDV, ensuring transparency and reproducibility.
TABLE I: A summary of various classification datasets used in our experiments.
Dataset Sample Input Number of Minor Class Data
Size Dimension Classes Proportion Type
2dplanes [64] 40768 10 2 0.499 Tabular
electricity [65] 38474 6 2 0.5 Tabular
fired [64] 40768 10 2 0.498 Tabular
BBC [66] 2225 768 5 0.17 Text
IMDB [67] 50000 768 2 0.5 Text
STL10 [68] 5000 96 10 0.01 Image
CIFAR10 [69] 50000 2048 10 0.1 Image
6.2 Runtime Efficiency Analysis
We assess the computational efficiency of NDDV against leading data valuation methods by measuring the time required to evaluate 1,0001{,}000 data samples on six representative datasets. This benchmark ensures both statistical reliability and manageable computational cost. As shown in Fig. 6(a), NDDV achieves the shortest runtime across all datasets while preserving high valuation accuracy.
Traditional combinatorial methods such as DataShapley and BetaShapley incur the highest runtime overhead due to exponential sample permutations, making them impractical for large-scale learning. Approximations like InfluenceFunction, DataBanzhaf, and LOO reduce computation time but remain slower than modern surrogates such as AME, KNNShapley, and Data-OOB. Although KNNShapley benefits from a closed-form estimator, its cost increases sharply with higher data dimensionality, and Data-OOB suffers from repetitive model retraining.
To evaluate scalability, we extend experiments to synthetic datasets of increasing size and feature dimensionality, with $n\in\{10^{4},10^{5},10^{6}\}$ and $d\in\{5,50,500\}$. As illustrated in Fig. 6(b), NDDV sustains near-linear growth in runtime, remaining efficient even at $(n,d)=(10^{6},500)$, where it completes valuation over 58× faster than the best-performing baseline. These results confirm that NDDV provides both scalability and practicality for large-scale data valuation.
6.3 Effectiveness Comparison
We assess the effectiveness of competing data valuation methods in detecting corrupted samples using the F1-score, which balances precision and recall in binary classification. The results in Table II report F1-scores for nine methods across six datasets under two training sizes (n=1,000n=1{,}000 and n=10,000n=10{,}000).
NDDV consistently achieves the best or second-best performance across all datasets and scales. When n=10,000n=10{,}000, its advantage becomes more pronounced, especially on complex benchmarks such as CIFAR10 and STL10, where dynamic valuation modeling yields substantial accuracy gains.
Among baselines, KNNShapley and Data-OOB perform competitively due to efficient approximations, yet both decline as dataset size increases. In contrast, LOO, AME, and InfluenceFunction perform poorly under large-scale settings, reflecting their limited capacity for robust mislabeled-data detection. DataShapley and BetaShapley improve marginally but remain unstable due to their combinatorial complexity.
Overall, the F1-score analysis demonstrates that NDDV delivers state-of-the-art effectiveness across all datasets and training conditions.
Figure 6: Elapsed time comparison between NDDV and existing methods.
TABLE II: F1-score of existing methods on the following datasets when (left) $n=1,000$ and (right) $n=10,000$.
Dataset $n=1,000$ $n=10,000$
LOO Data Beta Data Influence KNN AME Data NDDV KNN AME Data NDDV
Shapley Shapley Banzhaf Function Shapley -OOB Shapley -OOB
2dplanes $0.18\pm$ $0.17\pm$ $\pm 0.16±0.16\pm$ $0.18\pm$ $0.30\pm$ $0.18\pm$ $0.46\pm$ $\mathbf{0.67}\pm$ $37\pm 0.0$$01\pm$ $0.71\pm$ $\mathbf{0.79}\pm$
$0.003$ $0.005$ $0.003$ $0.009$ $0.005$ $0.007$ $0.009$ $0.007$ $0.005$ $0.004$ $0.012$ $0.002$ $0.005$
electricity $0.18\pm$ $0.17\pm$ $0.19\pm$ $0.18\pm$ $0.19\pm$ $23\pm 0.0$$01\pm$ $\mathbf{0.37}\pm$ $\underline{\emph{0.36}}\pm$ $32\pm 0.0$$01\pm$ $0.38\pm$ $\mathbf{0.44}\pm$
$0.004$ $0.004$ $0.006$ $0.002$ $0.003$ $0.006$ $0.010$ $0.002$ $0.002$ $0.001$ $0.009$ $0.003$ $0.002$
bbc $2\pm 0.11±0$$1\pm$ $0.11\pm$ $0.18\pm$ $0.16\pm$ $0.31\pm$ $0.11\pm$ $0.18\pm$ $\mathbf{0.86}\pm$ $52\pm 0.0$$01\pm$ $0.73\pm$ $\mathbf{0.85}\pm$
$0.004$ $0.004$ $0.003$ $0.005$ $0.002$ $0.008$ $0.009$ $0.004$ $0.002$ $0.005$ $0.010$ $0.002$ $0.006$
IMDB $12\pm 0.0$$9\pm 0.0$$09\pm$ $0.15\pm$ $0.16\pm$ $0.22\pm$ $0.18\pm$ $0.17\pm$ $\mathbf{0.27}\pm$ $0.29\pm$ $0.18\pm$ $0.48\pm$ $\mathbf{0.52}\pm$
$0.002$ $0.004$ $0.003$ $0.002$ $0.009$ $0.008$ $0.011$ $0.005$ $0.007$ $0.002$ $0.012$ $0.002$ $0.003$
STL10 $0.13\pm$ $0.17\pm$ $0.16\pm$ $0.18\pm$ $0.14\pm$ $28\pm 0.0$$01\pm$ $0.22\pm$ $\mathbf{0.71}\pm$ $16\pm 0.0$$01\pm$ $0.77$ $\mathbf{0.91}\pm$
$0.006$ $0.004$ $0.002$ $0.005$ $0.009$ $0.007$ $0.009$ $0.003$ $0.008$ $0.009$ $0.012$ $0.002$ $0.003$
CIFAR10 $0.18\pm$ $0.19\pm$ $0.20\pm$ $0.17\pm$ $0.19\pm$ $24\pm 0.0$$02\pm$ $0.40\pm$ $\mathbf{0.59}\pm$ $27\pm 0.0$$01\pm$ $0.46\pm$ $\mathbf{0.58}\pm$
$0.004$ $0.003$ $0.005$ $0.002$ $0.007$ $0.004$ $0.008$ $0.004$ $0.004$ $0.009$ $0.010$ $0.001$ $0.004$
  • Note: The mean and standard deviation of the F1-score are derived from 5 independent experiments. The highest and second-highest results are highlighted in bold and underlined, respectively.
6.4 Robustness Against Data Corruption and Manipulation
Real-world datasets often include mislabeled or noisy samples that impair model generalization. An effective data valuation method should downweight such corrupted points and accurately rank samples by their true contribution to learning.
Corrupted Data Detection. We inject controlled label and feature noise and measure detection accuracy using F1-scores in Fig. 7 (first column). NDDV consistently achieves the highest alignment with ground-truth corruption labels, outperforming strong baselines such as KNNShapley, AME, Data-OOB, and InfluenceFunction.
High-Value Data Removal. Following [5, 14], we iteratively remove the most valuable samples and re-train the model (Fig. 7, second column). NDDV causes the steepest test accuracy decline, confirming its precise identification of influential data, while AME and KNNShapley show weaker prioritization.
Low-Value Data Addition. Adding data from lowest to highest estimated value (Fig. 7, third column) further highlights robustness: NDDV matches or surpasses Data-OOB across most datasets, maintaining stable degradation patterns under noisy inclusion.
Performance under Noise. When label or feature noise increases from 5% to 45% (Fig. 9), NDDV maintains high detection stability, whereas baselines—especially AME and DataShapley—show large fluctuations. These findings confirm NDDV’s resilience and reliability for data filtering and quality-aware learning.
Noise Rate LOO Data Beta Data Influence KNN AME Data NDDV
Shapley Shapley Banzhaf Function Shapley -OOB
5% $0.09\pm$ $2\pm 0.11±0$$1\pm$ $0.09\pm$ $0.11\pm$ $17\pm 0.0$$01\pm$ $0.62\pm$ $\mathbf{0.74}\pm$
$0.003$ $0.007$ $0.008$ $0.004$ $0.003$ $0.003$ $0.009$ $0.002$ $0.003$
10% $0.16\pm$ $\pm 0.19±0.19\pm$ $\pm 0.18±0.18\pm$ $0.30\pm$ $0.18\pm$ $0.74\pm$ $\mathbf{0.76}\pm$
$0.007$ $0.010$ $0.009$ $0.005$ $0.003$ $0.003$ $0.010$ $0.002$ $0.003$
20% $0.30\pm$ $\pm 0.25±0.25\pm$ $\pm 0.31±0.31\pm$ $45\pm 0.0$$010\pm$ $\mathbf{0.79}\pm$ $\underline{\emph{0.77}}\pm$
$0.005$ $0.008$ $0.008$ $0.002$ $0.002$ $0.004$ $0.009$ $0.001$ $0.001$
30% $0.39\pm$ $0.52\pm$ $0.51\pm$ $\pm 0.42±0.42\pm$ $0.55\pm$ $0.46\pm$ $\mathbf{0.80}\pm$ $\underline{\emph{0.78}}\pm$
$0.003$ $0.012$ $0.010$ $0.002$ $0.008$ $0.002$ $0.011$ $0.001$ $0.004$
40% $4\pm 0.55±0$$5\pm$ $0.56\pm$ $0.48\pm$ $0.46\pm$ $0.60\pm$ $0.58\pm$ $0.73\pm$ $\mathbf{0.74}\pm$
$0.008$ $0.008$ $0.008$ $0.003$ $0.004$ $0.002$ $0.010$ $0.001$ $0.002$
45% $\pm 0.55±0.55\pm$ $0.62\pm$ $\pm 0.48±0.48\pm$ $0.56\pm$ $0.27\pm$ $0.63\pm$ $\mathbf{0.67}\pm$
$0.007$ $0.008$ $0.009$ $0.003$ $0.001$ $0.004$ $0.009$ $0.001$ $0.004$
  • Note: The mean and standard deviation of the F1-score are derived from 5 independent experiments. The highest and second-highest results are highlighted in bold and underlined, respectively.
TABLE III: F1-score of existing methods on the different label noise rates.
Noise Rate LOO Data Beta Data Influence KNN AME Data NDDV
Shapley Shapley Banzhaf Function Shapley -OOB
5% $0.09\pm$ $\pm 0.10±0.10\pm$ $0.07\pm$ $0.10\pm$ $17\pm 0.0$$09\pm$ $0.15\pm$ $\mathbf{0.30}\pm$
$0.007$ $0.009$ $0.007$ $0.004$ $0.003$ $0.003$ $0.012$ $0.002$ $0.006$
10% $\pm 0.18±0.18\pm$ $0.18\pm$ $\pm 0.15±0.15\pm$ $0.15\pm$ $0.18\pm$ $0.21\pm$ $\mathbf{0.28}\pm$
$0.007$ $0.010$ $0.009$ $0.005$ $0.003$ $0.003$ $0.010$ $0.002$ $0.003$
20% $33\pm 0.0$$1\pm 0.0$$01\pm$ $0.28\pm$ $0.30\pm$ $27\pm 0.0$$01\pm$ $0.32\pm$ $\mathbf{0.34}\pm$
$0.005$ $0.008$ $0.008$ $0.002$ $0.002$ $0.002$ $0.010$ $0.001$ $0.003$
30% $43\pm 0.0$$1\pm 0.0$$01\pm$ $0.33\pm$ $\pm 0.35±0.35\pm$ $0.01\pm$ $0.37\pm$ $\mathbf{0.45}\pm$
$0.008$ $0.012$ $0.010$ $0.002$ $0.008$ $0.002$ $0.012$ $0.001$ $0.005$
40% $51\pm 0.0$$1\pm 0.0$$1\pm 0.0$$01\pm$ $0.37\pm$ $40\pm 0.0$$01\pm$ $0.43\pm$ $\mathbf{0.57}\pm$
$0.008$ $0.010$ $0.008$ $0.003$ $0.004$ $0.002$ $0.010$ $0.001$ $0.004$
45% $53\pm 0.0$$1\pm 0.0$$01\pm$ $0.50\pm$ $0.47\pm$ $39\pm 0.0$$01\pm$ $0.46\pm$ $\mathbf{0.62}\pm$
$0.007$ $0.011$ $0.009$ $0.003$ $0.001$ $0.002$ $0.012$ $0.001$ $0.006$
  • Note: The mean and standard deviation of the F1-score are derived from 5 independent experiments. The highest and second-highest results are highlighted in bold and underlined, respectively.
TABLE IV: F1-score of existing methods on the different feature noise rates.
Figure 7: Data valuation experiments on the following datasets with 10% label noisy rate. (First column) Detecting corrupted data experiment. (Second column) Removing high-value data experiment. (Third column) Adding low-value data experiment.
Figure 8: Noisy data detection task on the following datasets. a. Noisy label data detection task. b. Noisy feature data detection task. The F1-score of various methods is compared on the six noise proportion settings. The higher F1-score indicates superior performance.
6.5 Ablation Study
The ablation study was conducted to examine how each component of NDDV contributes to its overall performance, fairness, and stability. Figure V presents the comparative results under various configurations, where individual modules were selectively removed or replaced.
When the mean-field controller was disabled, both Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD) increased significantly, showing that fairness improvement in NDDV arises directly from the adaptive mean-field interactions. Removing the dynamic state encoder led to a noticeable drop in F1-score and rank stability, as the model could no longer capture temporal dependencies between evolving data representations. Similarly, substituting the interpretable neural valuator with a standard multilayer perceptron degraded Rank Consistency (RC) and Spearman Correlation (SC) by over 20%, validating that the Kolmogorov–Arnold representation [46, 47] enhances the consistency and interpretability of valuation.
These findings demonstrate that all three components are indispensable. The dynamic encoder ensures robust representation of sample influence, the mean-field controller guarantees fairness-aware value propagation, and the neural valuator provides interpretability without compromising efficiency. Together, they form a cohesive control-driven framework that balances accuracy, equity, and transparency.
TABLE V: Ablation study of NDDV components on the average performance across six datasets. Each component is removed or replaced to examine its contribution to effectiveness, fairness, and interpretability. Bold values indicate the best results.
Configuration F1 (%) SPD ($\downarrow$) RC (%) SC (%)
Full NDDV (All modules) 92.4 0.03 95.7 94.8
w/o Dynamic State Encoder (DSE) 85.2 0.07 78.9 80.1
w/o Mean-Field Controller (MFC) 88.1 0.11 90.3 91.4
w/o Interpretable Neural Valuator (INV) 89.6 0.05 74.8 73.2
6.6 Fairness Evaluation
This experiment investigates how effectively NDDV mitigates valuation bias and ensures equitable contribution assessment across heterogeneous data groups. The study analyzes feature bias and label noise bias following the established evaluation design in fairness-aware machine learning [62, 70, 71, 72]. The quantitative results are visualized in Figure 9.
In the feature-bias test, the data distribution was intentionally skewed by introducing correlations between sensitive attributes (e.g., gender or age) and predictive features. Under this setting, traditional valuation approaches—such as InfluenceFunction [31] and DataShapley [5]—exhibited amplified bias, reflected by high Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD) values. In contrast, NDDV maintained near-zero parity gaps due to its mean-field controller, which regularizes global data interactions through adaptive reweighting.
In the label-noise scenario, 15% of the labels were randomly flipped to simulate annotation errors. Here, Data-OOB [14] and AME [13] partially reduced bias but suffered from unstable group-level valuations. NDDV achieved the lowest SPD and EOD across all noise levels, demonstrating its robustness to both systematic and random unfairness sources. The fairness improvement arises from the dynamic correction mechanism, which adjusts each sample’s influence as the control process evolves, ensuring balanced value propagation even when the dataset becomes corrupted.
These results confirm that fairness constraints embedded in the stochastic control formulation enable NDDV to allocate value equitably without compromising model performance. The consistent superiority observed in Figure 9 indicates that dynamic mean-field adjustment is an effective principle for mitigating data bias in valuation tasks.
Figure 9: Fairness evaluation on the following datasets. a. Fairness evaluation with the EOp-score. The lower EOp-score and the higher F1-score indicate superior performance. b. Fairness evaluation with the EOdds-score. The lower EOdds-score and the higher F1-score indicate superior performance.
6.7 Interpretability Analysis
To evaluate the transparency and reliability of the proposed valuation process, Figure 10 visualizes the temporal evolution of data values during training under different levels of label noise. Each curve represents how the estimated contribution of a data point changes as the model learns, providing insights into the internal learning dynamics of NDDV.
The visual patterns in Figure 10 show that high-value samples maintain smooth and monotonic growth trajectories, signifying their stable and consistent contribution to improving model generalization. In contrast, noisy or mislabeled samples exhibit oscillating or declining valuation curves, reflecting unstable gradients and low utility to the learning process. These results demonstrate that NDDV inherently distinguishes informative samples from detrimental ones, even when trained under corrupted supervision.
This interpretability arises from the stochastic control formulation of NDDV, which traces each sample’s influence through its dynamic state and control gradients. The learned trajectories offer a clear, human-understandable interpretation of why a data point gains or loses importance over time. Such temporal explainability not only improves transparency but also supports practical auditing, enabling users to identify mislabeled data and assess data quality at both individual and population levels.
Through this analysis, NDDV provides an interpretable and trustworthy framework that integrates valuation, fairness, and data quality understanding within a single unified system.
Figure 10: Noisy data detection task for the interpretable NDDV on the following datasets. a. Noisy label data detection task. b. Noisy feature data detection task. The F1-score of various methods is compared on the six noise proportion settings. The higher F1-score indicates superior performance.
6.8 Sensitivity Analysis
We conduct a sensitivity analysis to examine how the performance of NDDV varies with changes in key hyperparameters. All experiments are performed on the PLC dataset using the mislabeled data detection task. The results, summarized in Fig. 11, provide insights into the stability and robustness of NDDV under different configurations.
Effect of Data Re-weighting. We assess the impact of enabling or disabling the re-weighting of data points. As shown in Fig. 11a, including re-weighting significantly improves performance in mislabeled data detection and utility-based data manipulation. When re-weighting is disabled, the model exhibits a clear decline in detection accuracy, highlighting the importance of learning adaptive sample weights to emphasize informative or suspicious points.
Effect of Mean-Field Interaction Strength $a$. The hyperparameter $a$ controls the strength of mean-field interactions between data points. We test values $a\in\{1,3,5,10\}$ and report the corresponding detection performance in Fig. 11b. While performance remains stable for $a=1,3,5$, we observe degradation when $a=10$, suggesting that overly strong interactions may lead to over-smoothing or unwanted interference between samples.
Effect of Diffusion Constant $\sigma$. We evaluate NDDV’s sensitivity to the diffusion constant $\sigma$, which governs the randomness in the dynamic process. Fig. 11c shows results for $\sigma\in\{0.001,0.01,0.1,1.0\}$. The performance deteriorates markedly at $\sigma=1.0$, indicating that excessive noise disrupts the learning dynamics. Lower values yield stable and accurate results, confirming the need for moderate diffusion strength.
Effect of Meta-Data Size. We vary the size of the meta-dataset used for meta-learning, testing sizes {10, 100, 300}. As shown in Fig. 11d, the detection performance remains relatively stable across these settings. This indicates that NDDV does not require a large meta-dataset to achieve strong performance, making it practical for scenarios with limited labeled or auxiliary data.
Effect of Meta Hidden Layer Size. We analyze sensitivity to the hidden layer size in the meta-network. Fig. 11e reports results when the hidden size is reduced to 5. Performance drops significantly in this case, suggesting that insufficient model capacity can hinder the learning of meaningful valuation dynamics. Larger hidden sizes help avoid underfitting and improve the stability of utility estimation.
Figure 11: Sensitivity analysis for NDDV. Using the detecting corrupted data task as an example: a. Impact of data points re-weighting. b. Impact of the mean-field interactions. c. Impact of the diffusion constant. d. Impact of the metadata sizes. e. Impact of the meta hidden points.
Summary. Overall, NDDV demonstrates strong robustness to moderate changes in hyperparameter values. The method is particularly stable across meta-dataset sizes and interaction strengths, while maintaining high detection performance when diffusion and model capacity are appropriately calibrated.
7 Conclusion
This work introduced Neural Dynamic Data Valuation (NDDV), a framework that formulates data valuation as a stochastic optimal control process to address inefficiency, fairness imbalance, and interpretability limitations in existing methods. NDDV models valuation as continuous dynamics, where each data point interacts with a mean-field state to determine its evolving importance. The framework integrates three innovations: a one-pass stochastic control formulation eliminating repetitive retraining; a fairness-aware mean-field reweighting mechanism ensuring equitable contribution; and an interpretable Kolmogorov–Arnold network with Matérn kernels that reveals how data value evolves through training. Experiments across six datasets show that NDDV achieves superior F1-scores, enhanced fairness, and up to 58× runtime improvement over state-of-the-art methods, confirming its scalability and transparency.
Future work will extend NDDV to dynamic data markets, enabling real-time valuation under concept drift, and to federated multimodal settings, where privacy-preserving mechanisms are required for heterogeneous and distributed data environments.

### v5

@@
 1 Introduction
-Data has become a central asset in the modern economy, functioning as intellectual property that drives value creation in data marketplaces [1, 2, 3].
+Data has become a central asset in the modern economy, functioning as intellectual property that drives value creation in data marketplaces [agarwal2019, tian2022private, pei2020survey].
 The intrinsic worth of data facilitates sharing, exchange, and reuse across organizations, businesses, and research communities.
@@
 Earlier research on data valuation has evolved from model-specific influence estimation to cooperative game formulations and recent efficiency-driven approximations.
-The influence-based approach [4] estimates each sample’s importance through sensitivity analysis of model parameters but struggles to capture complex dependencies among data points.
-The game-theoretic framework [5, 6, 7, 8, 9] assigns value through average marginal contributions, offering strong fairness principles but requiring exponentially many evaluations.
-To mitigate this burden, efficient approximation techniques have been introduced, such as neighbor-based models [10, 11, 12] and ensemble reweighting strategies [13, 14], yet these often compromise fairness or interpretability.
-Further, explainability-oriented studies [15] enhance transparency in local model reasoning but do not extend to global data valuation.
+The influence-based approach [koh2017understanding] estimates each sample’s importance through sensitivity analysis of model parameters but struggles to capture complex dependencies among data points.
+The game-theoretic framework [ghorbani2019, shapley1953value, ghor2020dis, schoch2022cs, kwon2022beta] assigns value through average marginal contributions, offering strong fairness principles but requiring exponentially many evaluations.
+To mitigate this burden, efficient approximation techniques have been introduced, such as neighbor-based models [jia2019, jia2019b, kwon2021efficient] and ensemble reweighting strategies [lin2022measuring, kwon2023dataoob], yet these often compromise fairness or interpretability.
+Further, explainability-oriented studies [chen2022explaining] enhance transparency in local model reasoning but do not extend to global data valuation.
 Despite these advances, existing methods remain limited by three unresolved issues: excessive computational cost that restricts scalability, biased or inconsistent value allocation across heterogeneous data, and lack of transparency regarding how value evolves during learning.
-These unresolved problems motivate a unified framework that models data value dynamically, enabling efficient, fair, and interpretable estimation at scale [1, 2, 3].
+These unresolved problems motivate a unified framework that models data value dynamically, enabling efficient, fair, and interpretable estimation at scale [agarwal2019, tian2022private, pei2020survey].
 To overcome these limitations, we propose Neural Dynamic Data Valuation (NDDV), which redefines data valuation as a stochastic optimal control problem.
@@
 The dynamical interpretation of deep neural networks (DNNs) as continuous-time systems has introduced new theoretical perspectives on learning and optimization.
-Foundational studies view deep residual architectures as discrete approximations of ordinary differential equations (ODEs), bridging neural network propagation and numerical integration theory [16, 17, 18].
-This connection enables analysis in the continuum limit through Wasserstein geometry, which links gradient flow and optimal transport [19].
-The resulting framework naturally extends to mean-field control formulations for learning dynamics [20, 21].
-Building on this perspective, several optimization algorithms have been reformulated as control problems [22, 23].
-Stochastic gradient descent (SGD) has been analyzed through the lens of stochastic dynamical systems, where Langevin and Lévy dynamics describe implicit regularization effects [24, 25].
-Further, studies of Gram matrix evolution have demonstrated convergence and global optimality properties [26, 27], providing guidance for adaptive tuning of hyperparameters such as learning rate and batch size [28, 29].
+Foundational studies view deep residual architectures as discrete approximations of ordinary differential equations (ODEs), bridging neural network propagation and numerical integration theory [weinan2017proposal, he2016deep, lu2017beyond].
+This connection enables analysis in the continuum limit through Wasserstein geometry, which links gradient flow and optimal transport [JMLR:v20:16-243].
+The resulting framework naturally extends to mean-field control formulations for learning dynamics [hu2017control, han2018mean].
+Building on this perspective, several optimization algorithms have been reformulated as control problems [li2017maximum, li2018optimal].
+Stochastic gradient descent (SGD) has been analyzed through the lens of stochastic dynamical systems, where Langevin and Lévy dynamics describe implicit regularization effects [pavliotis2014stochastic, simsekli2019tail].
+Further, studies of Gram matrix evolution have demonstrated convergence and global optimality properties [du2018gradienta, du2018gradientb], providing guidance for adaptive tuning of hyperparameters such as learning rate and batch size [li2017stochastic, an2018stochastic].
 Collectively, these results suggest that the theory of dynamic systems and optimal control offers a principled foundation for reinterpreting neural learning and motivates the control-based formulation adopted in this work.
@@
 Data valuation has emerged as a key research topic to quantify each sample’s contribution to model performance.
-Early approaches such as Leave-One-Out (LOO) [4] and Data Shapley [10, 5] estimate marginal contributions by iteratively removing data points and observing model performance changes.
+Early approaches such as Leave-One-Out (LOO) [koh2017understanding] and Data Shapley [jia2019, ghorbani2019] estimate marginal contributions by iteratively removing data points and observing model performance changes.
 Although theoretically grounded, these methods suffer from exponential computational cost.
-Extensions like Beta Shapley [9] and Data Banzhaf [30] improve fairness modeling but remain computationally demanding.
-Influence-based estimators [31] leverage gradient approximations to reduce complexity, yet often lose robustness in nonconvex learning settings.
+Extensions like Beta Shapley [kwon2022beta] and Data Banzhaf [wang2022data] improve fairness modeling but remain computationally demanding.
+Influence-based estimators [feldman2020neural] leverage gradient approximations to reduce complexity, yet often lose robustness in nonconvex learning settings.
 To enhance scalability, proxy and ensemble-based techniques were introduced.
-KNNShapley [11] employs a $k$-Nearest Neighbor classifier as a learning-agnostic surrogate, while AME [13] and Data-OOB [14] adopt ensemble reweighting and out-of-bag estimators for efficiency.
+KNNShapley [jia2019b] employs a $k$-Nearest Neighbor classifier as a learning-agnostic surrogate, while AME [lin2022measuring] and Data-OOB [kwon2023dataoob] adopt ensemble reweighting and out-of-bag estimators for efficiency.
 These methods, however, still struggle to maintain fairness and transparency.
-Algorithm-agnostic formulations, such as volume-based valuation [32] and Wasserstein-distance metrics [33], bypass model retraining but lack the ability to detect label errors or explain valuation outcomes.
-Beyond standard utility estimation, marginal contribution analysis has been extended to feature attribution [34, 35], interpretability [36, 37], and collaborative learning [38, 39].
-Further refinements [40, 41] relax classical Shapley axioms to improve efficiency, while reinforcement learning-based estimators [42] model data usage likelihood during training.
+Algorithm-agnostic formulations, such as volume-based valuation [xu2021validation] and Wasserstein-distance metrics [just2023lava], bypass model retraining but lack the ability to detect label errors or explain valuation outcomes.
+Beyond standard utility estimation, marginal contribution analysis has been extended to feature attribution [lundberg2017unified, covert2021explaining], interpretability [stier2018analysing, ghorbani2020neuron], and collaborative learning [sim2020collaborative, xu2021gradient].
+Further refinements [yan2021if, rozemberczki2022shapley] relax classical Shapley axioms to improve efficiency, while reinforcement learning-based estimators [yoon2020data] model data usage likelihood during training.
 Despite these developments, current frameworks still face three inherent constraints: high computational cost, unfair valuation under data heterogeneity, and limited interpretability.
@@
 which quantifies the change in utility when the $i$-th point is excluded from the full dataset.
-This metric approximates the influence of a data point but can yield inaccurate values close to zero in practice [43].
+This metric approximates the influence of a data point but can yield inaccurate values close to zero in practice [basu2020influence].
 Definition 2 (Static Marginal Contribution).
@@
 NDDV comprises three tightly integrated modules that operate coherently to achieve these objectives.
-The first module, the Dynamic State Encoder, transforms data samples into a latent representation governed by stochastic differential equations, which model their dynamic interactions with the evolving learning system [24, 23].
-The second module, the Fairness-Aware Mean-Field Controller, defines a mean-field potential that regulates global fairness across samples by adjusting their control weights according to their statistical influence, ensuring equitable value allocation even under data heterogeneity [44, 45].
-The third module, the Interpretable Neural Valuator, leverages KANs with Matérn kernels [46, 47] to map the evolving latent states into valuation scores, preserving the underlying dynamics while offering transparency in the contribution estimation.
+The first module, the Dynamic State Encoder, transforms data samples into a latent representation governed by stochastic differential equations, which model their dynamic interactions with the evolving learning system [pavliotis2014stochastic, li2018optimal].
+The second module, the Fairness-Aware Mean-Field Controller, defines a mean-field potential that regulates global fairness across samples by adjusting their control weights according to their statistical influence, ensuring equitable value allocation even under data heterogeneity [carmona2013control, bensoussan2017linear].
+The third module, the Interpretable Neural Valuator, leverages KANs with Matérn kernels [liu2024kan, rasmussen2003gaussian] to map the evolving latent states into valuation scores, preserving the underlying dynamics while offering transparency in the contribution estimation.
 Collectively, these modules constitute a unified control-driven valuation framework that jointly ensures efficiency, fairness and interpretability.
@@
 4.1 Learning Dynamic Valuation Trajectories via Stochastic Control
-To address the inefficiency and static limitations of conventional data valuation, we formulate valuation as a stochastic optimal control problem [48, 49].
+To address the inefficiency and static limitations of conventional data valuation, we formulate valuation as a stochastic optimal control problem [casert2024learning, gao2024learning].
 This dynamic view models how data points jointly influence learning over time, where each point follows a stochastic trajectory capturing its evolving contribution (see Fig.
@@
 where $b$ is the drift function, $\sigma$ is the diffusion coefficient, and $W_{t}$ is a standard Wiener process.
-Applying the Stochastic Maximum Principle (SMP) [50], we introduce the Hamiltonian:
+Applying the Stochastic Maximum Principle (SMP) [kushner1965stochastic], we introduce the Hamiltonian:
 $\mathcal{H}(X_{t},Y_{t},Z_{t},\psi_{t})=b(X_{t},\psi_{t})\!\cdot\!Y_{t}+\mathrm{tr}(\sigma^{\top}Z_{t})-R(X_{t},\psi_{t}),$ (6)
@@
 $\mathrm{d}Y_{t}=-\nabla_{x}\mathcal{H}(X_{t},Y_{t},\psi_{t})\,\mathrm{d}t+Z_{t}\,\mathrm{d}W_{t},\qquad Y_{T}=-\nabla_{x}\Phi(X_{T}).$ (7)
-The coupled FSDE–BSDE system forms a McKean–Vlasov FBSDE [51, 44, 52], and the optimal control $\psi_{t}^{*}$ satisfies:
+The coupled FSDE–BSDE system forms a McKean–Vlasov FBSDE [carmona2015forward, carmona2013control, sen2016nonlinear], and the optimal control $\psi_{t}^{*}$ satisfies:
 $\mathcal{H}(X_{t}^{*},Y_{t}^{*},Z_{t}^{*},\psi_{t}^{*})\geq\mathcal{H}(X_{t}^{*},Y_{t}^{*},Z_{t}^{*},\psi),\quad\forall\,\psi.$ (8)
@@
 $=-X_{i,T}\!\cdot\!Y_{i,T},$ (9)
-linking valuation to the gradient of the cost function [53, 54].
+linking valuation to the gradient of the cost function [shrikumar2017learning, ancona2019explaining].
 The marginal and dynamic valuation score is defined as:
@@
 allowing continuous tracking of sample importance during training.
-This stochastic dynamic formulation eliminates combinatorial retraining, models temporal evolution of data influence, and offers interpretability across layers and epochs, the capabilities absent in static valuation heuristics [55, 56].
+This stochastic dynamic formulation eliminates combinatorial retraining, models temporal evolution of data influence, and offers interpretability across layers and epochs, the capabilities absent in static valuation heuristics [serban_cvodes_2005, jorgensen_2007].
 Figure 2: Learning data stochastic dynamic schematic.
@@
 Adaptive Reweighting via Meta-Learned Weights.
-Each data point is dynamically reweighted by a meta-learned function $V(\Phi_{i}(X_{T});\theta)$ inspired by Stackelberg game theory [45].
+Each data point is dynamically reweighted by a meta-learned function $V(\Phi_{i}(X_{T});\theta)$ inspired by Stackelberg game theory [bensoussan2017linear].
 The control objective becomes:
@@
 $\begin{cases}\psi^{*}(\theta)=\arg\min_{\psi}L(\psi;\theta),\\ \theta^{*}=\arg\min_{\theta}\frac{1}{M}\sum_{i=1}^{M}\ell_{i}(\psi^{*}(\theta)),\end{cases}$ (13)
-solved efficiently through a single-loop online update [57, 58]. This procedure amplifies representative and fair samples while reducing bias from noisy or minority points.
-Mean-Field Interaction and Weighted Aggregation. Group fairness is enforced by embedding reweighted data into a mean-field dynamic system [44]:
+solved efficiently through a single-loop online update [shu2019meta, yang2024curriculum]. This procedure amplifies representative and fair samples while reducing bias from noisy or minority points.
+Mean-Field Interaction and Weighted Aggregation. Group fairness is enforced by embedding reweighted data into a mean-field dynamic system [carmona2013control]:
 $\mathrm{d}X_{i,t}=b(X_{i,t},\mu_{t},\psi_{i,t})\,\mathrm{d}t+\sigma\,\mathrm{d}W_{i,t},$ (14)
@@
 $\mu_{t}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{V}(\Phi_{i}(\cdot,\psi_{i,T});\theta)X_{i,t}.$ (15)
-With a linear–quadratic drift [59, 45, 60],
+With a linear–quadratic drift [yong2013linear, bensoussan2017linear, carmona2015mean],
 $\mathrm{d}X_{t}=[a(\mu_{t}-X_{t})+\psi_{t}]\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t},$ (16)
@@
 Formal Fairness Guarantee.
-Following [5, 61], NDDV satisfies symmetry, monotonicity, and null-player properties.
-Fairness is quantified by the Equal Opportunity (EOp) and Equalized Odds (EOdds) scores [62]:
+Following [ghorbani2019, sim2022data], NDDV satisfies symmetry, monotonicity, and null-player properties.
+Fairness is quantified by the Equal Opportunity (EOp) and Equalized Odds (EOdds) scores [hardt2016equality]:
 $\begin{cases}\text{EOp}_{i}=|\text{TPR}-\text{TPR}_{i}|,\\ \text{EOdds}_{i}=|\text{TPR}-\text{TPR}_{i}|+|\text{FPR}-\text{FPR}_{i}|.\end{cases}$ (19)
@@
 The drift term $b(X_{t},\mu_{t},\psi_{t})$, which governs data evolution, depends on both the control policy $\psi$ and the reweighting parameters $\mathcal{V}$.
-Inspired by KANs [46], we replace dense weight matrices with univariate basis functions, expressing $\psi$ and $\mathcal{V}$ as compositions of nonlinear mappings for improved interpretability and analytical control.
+Inspired by KANs [liu2024kan], we replace dense weight matrices with univariate basis functions, expressing $\psi$ and $\mathcal{V}$ as compositions of nonlinear mappings for improved interpretability and analytical control.
 The control strategy $\psi$ is represented as:
@@
 where $h_{b}$ uses the SiLU function, and $h_{k}$ is a radial basis function (RBF) centered at $c$. Coefficients $\alpha_{b}$ and $\alpha_{k}$ are initialized via Xavier and optimized during training.
-Because Gaussian RBFs can over-smooth noisy distributions, we adopt a **Matérn kernel** [47] to improve local adaptivity:
+Because Gaussian RBFs can over-smooth noisy distributions, we adopt a **Matérn kernel** [rasmussen2003gaussian] to improve local adaptivity:
 $h_{k}(X_{t},c)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\,|X_{t}-c|}{\ell}\right)^{\!\nu}K_{\nu}\!\left(\frac{\sqrt{2\nu}\,|X_{t}-c|}{\ell}\right),$
@@
 Datasets.
-We evaluate our approach using six publicly available datasets provided in OpenDataVal [63], several of which have been widely adopted in prior work [5, 9].
+We evaluate our approach using six publicly available datasets provided in OpenDataVal [jiang2023opendataval], several of which have been widely adopted in prior work [ghorbani2019, kwon2022beta].
 These datasets span diverse domains—including image, text, and tabular data with varying degrees of class imbalance, feature sparsity, and task complexity, as detailed in Table I.
@@
 Baselines.
-We compare NDDV against eight representative data valuation methods: LOO [4], DataShapley [5], BetaShapley [9], DataBanzhaf [30], InfluenceFunction [31], KNNShapley [11], AME [13], and Data-OOB [14].
+We compare NDDV against eight representative data valuation methods: LOO [koh2017understanding], DataShapley [ghorbani2019], BetaShapley [kwon2022beta], DataBanzhaf [wang2022data], InfluenceFunction [feldman2020neural], KNNShapley [jia2019b], AME [lin2022measuring], and Data-OOB [kwon2023dataoob].
 These methods cover a wide spectrum of valuation paradigms, including exact and approximate game-theoretic methods, influence-based techniques, and ensemble-based approaches.
@@
 Size Dimension Classes Proportion Type
-2dplanes [64] 40768 10 2 0.499 Tabular
-electricity [65] 38474 6 2 0.5 Tabular
-fired [64] 40768 10 2 0.498 Tabular
-BBC [66] 2225 768 5 0.17 Text
-IMDB [67] 50000 768 2 0.5 Text
-STL10 [68] 5000 96 10 0.01 Image
-CIFAR10 [69] 50000 2048 10 0.1 Image
+2dplanes [feurer-arxiv19a] 40768 10 2 0.499 Tabular
+electricity [gama2004learning] 38474 6 2 0.5 Tabular
+fired [feurer-arxiv19a] 40768 10 2 0.498 Tabular
+BBC [greene2006practical] 2225 768 5 0.17 Text
+IMDB [maas2011learning] 50000 768 2 0.5 Text
+STL10 [coates2011analysis] 5000 96 10 0.01 Image
+CIFAR10 [krizhevsky2009learning] 50000 2048 10 0.1 Image
 6.2 Runtime Efficiency Analysis
@@
 High-Value Data Removal.
-Following [5, 14], we iteratively remove the most valuable samples and re-train the model (Fig.
+Following [ghorbani2019, kwon2023dataoob], we iteratively remove the most valuable samples and re-train the model (Fig.
 7, second column).
@@
 Removing the dynamic state encoder led to a noticeable drop in F1-score and rank stability, as the model could no longer capture temporal dependencies between evolving data representations.
-Similarly, substituting the interpretable neural valuator with a standard multilayer perceptron degraded Rank Consistency (RC) and Spearman Correlation (SC) by over 20%, validating that the Kolmogorov–Arnold representation [46, 47] enhances the consistency and interpretability of valuation.
+Similarly, substituting the interpretable neural valuator with a standard multilayer perceptron degraded Rank Consistency (RC) and Spearman Correlation (SC) by over 20%, validating that the Kolmogorov–Arnold representation [liu2024kan, rasmussen2003gaussian] enhances the consistency and interpretability of valuation.
 These findings demonstrate that all three components are indispensable.
@@
 This experiment investigates how effectively NDDV mitigates valuation bias and ensures equitable contribution assessment across heterogeneous data groups.
-The study analyzes feature bias and label noise bias following the established evaluation design in fairness-aware machine learning [62, 70, 71, 72].
+The study analyzes feature bias and label noise bias following the established evaluation design in fairness-aware machine learning [hardt2016equality, chouldechova2017fair, zafar2017fairness, berk2021fairness].
 The quantitative results are visualized in Figure 9.
 In the feature-bias test, the data distribution was intentionally skewed by introducing correlations between sensitive attributes (e.g., gender or age) and predictive features.
-Under this setting, traditional valuation approaches—such as InfluenceFunction [31] and DataShapley [5]—exhibited amplified bias, reflected by high Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD) values.
+Under this setting, traditional valuation approaches—such as InfluenceFunction [feldman2020neural] and DataShapley [ghorbani2019]—exhibited amplified bias, reflected by high Statistical Parity Difference (SPD) and Equal Opportunity Difference (EOD) values.
 In contrast, NDDV maintained near-zero parity gaps due to its mean-field controller, which regularizes global data interactions through adaptive reweighting.
 In the label-noise scenario, 15% of the labels were randomly flipped to simulate annotation errors.
-Here, Data-OOB [14] and AME [13] partially reduced bias but suffered from unstable group-level valuations.
+Here, Data-OOB [kwon2023dataoob] and AME [lin2022measuring] partially reduced bias but suffered from unstable group-level valuations.
 NDDV achieved the lowest SPD and EOD across all noise levels, demonstrating its robustness to both systematic and random unfairness sources.
@@
 Future work will extend NDDV to dynamic data markets, enabling real-time valuation under concept drift, and to federated multimodal settings, where privacy-preserving mechanisms are required for heterogeneous and distributed data environments.
+[Uncaptioned image] Zhangyong Liang is working towards the PhD degree at the National Center for Applied Mathematics, Tianjin University.
+His research interests include AI for science, physics-informed machine learning, neural PDE solvers, data integration, and explainability.
+[Uncaptioned image] Ji Zhang is currently a full Professor in Computer Science at the University of Southern Queensland (UniSQ), Australia.
+He is an RSA Life Fellow, IET Fellow, BCS Fellow, IAAST Fellow, IEEE Senior Member, Australian Endeavour Fellow, Queensland International Fellow (Australia), and Izaak Walton Killam Scholar (Canada).
+His research interests span big data analytics, data intelligence, intelligent computing, computational intelligence and data privacy.
+He has published over 380 papers in major peer-reviewed international journals and conferences including TPAMI, TKDE, TKDD, TDSC, TIFS, TIST, CSUR, AAAI, IJCAI, VLDB, CIKM, SIGKDD, ICDE, ICDM, CVPR and WWW.
+[Uncaptioned image] Xin Wang received the Ph.D.
+degree from the University of Edinburgh in 2013.
+His research mainly focuses on Big Data, Knowledge Engineering and Natural Language Processing.
+He is a professor in the School of Computer Science and software engineering, Southwest Petroleum University, and serves as a Sichuan Provincial Distinguished Expert, China.
+He has published papers in the prestigious conferences and journals, including SIGMOD, VLDB, ICDE, CVPR and the IEEE Transactions on Knowledge and Data Engineering, Expert Systems with Applications, and received ICDE 2014 Best Paper Runner-up Award, WISE 2019 Best Paper Runner-up Award, etc.
+[Uncaptioned image] Pengfei Zhang received a Ph.D from Beijing University of Posts and Telecommunications in China, with a major in Computer Science.
+He has served as the reviewer, meta-reviewer or area chair for many top-tier conferences and journals.
+He has published more than 30 papers in NeurIPS, TKDE, IJCAI, TIFS, AAAI, CIKM and so on.
+He is also a guest editor for several journals.
+His current research interest focuses on big data, privacy protection and mobile crowdsensing.
+[Uncaptioned image] Zhao Li received his PhD degree in Computer Science from the University of Vermont in 2012, receiving the Excellent Graduate Award.
+He is currently an Adjunct Professor at Zhejiang University and co-founder of Hangzhou Yugu Technology Co., Ltd.
+He has published numerous papers in prestigious conferences and journals, including WWW, AAAI, IAAI, ICDE, VLDB, and TKDE.
+His research interests include machine learning, big data-driven applications, and large-scale graph computing.